A trader depreciates his vehicles at 40% per annum using the reducing balance method, with time proportionate depreciation in the year of acquisition. In addition to vehicles acquired on 1 July 2010 for R180 000, he acquired a vehicle on 1 May 2011 for R60 000. His depreciation for year ended 31.12.2011 will be_____.

Select one:

a. R81 600

b. R73 600

c. R83 200

d. R59 200


1
Expert's answer
2021-05-26T10:30:12-0400
"180000\times 0.4\times\frac{6}{12}=144000"

"144000\times0.4=57600"

"60000\times0.4\times\frac{8}{12}=16000"

"57600+16000=73600"

b. R73 600